# LibEER: A Comprehensive Benchmark and Algorithm Library for EEG-based Emotion Recognition

### v2

@@ III-D Datasets
Research in the EER field exhibits variability in dataset selection and often lacks clarity regarding the specific data segments utilized. This study evaluates the performance of popular models using four datasets, which are selected by comprehensively considering factors such as their frequency of use, data scale, and data quality. They will be introduced along with the relevant data segments, as summarized in Table III.
@@ III-D3 DEAP
The DEAP[22] dataset consists of 32 subjects (16 males, 16 females) who participated in one session containing 40 trials. Each trial involved a 1-minute music video designed to evoke specific emotions, followed by participant ratings for arousal, valence, dominance, and liking on a 1-9 scale. EEG data were recorded using a 32-channel cap at 512 Hz, and the dataset includes both raw and downsampled (128 Hz) versions. All data are used for both subject-dependent and cross-subject tasks.
@@ III-D4 MAHNOB-HCI
The MAHNOB-HCI[24] dataset includes 30 subjects (13 males, 17 females), though EEG data are completely missing for 2 subjects and partially missing for 3. Each subject participated in a single session with 20 trials, where videos lasting 34.9 to 117 seconds evoked emotions. After each video, participants rated arousal, valence, dominance, and predictability on a 1-9 scale. EEG data were recorded using a 32-channel system at 512 Hz, and the dataset includes both raw and downsampled (128 Hz) versions. Both complete and partial data are used for subject-dependent and cross-subject tasks.
@@ III-E Evaluation Methods and Metrics
where TP (True Positives), TN (True Negatives), FP (False Positives), and FN (False Negatives) are the metrics used to calculate accuracy and F1 score.
@@ IV LibEER Toolkit Framework
LibEER offers a user-friendly and equitable platform for EER implemented using PyTorch. It comprises three primary modules: the data loader, data splitting, and model training and evaluation. The data loader module standardizes the data formats across various datasets and provides a range of preprocessing techniques. The data splitting module delivers a cohesive solution applicable to different experimental tasks and partitioning methods. The model training and evaluation module presents a standardized and extensible framework for conducting model training and evaluation. Fig. 3 illustrates the framework of LibEER. Below, we provide detailed explanations and usage examples for each module.
Figure 4: The example shows how to LibEER to load data
Figure 5: The example how to use LibEER to slit data
@@ IV-1 Data loader
Additionally, LibEER provides the preprocess() function, which allows for the preprocessing of data based on user-defined settings. The preprocess() function operates on a trial basis and encompasses three primary steps: noise removal, feature extraction, and sample segmentation. When utilizing raw data directly from the dataset, only sample segmentation is executed. Noise removal aims to eliminate electromyography (EMG) noise, electrooculography (EOG) artifacts, and other forms of interference from the EEG signals. The feature extraction process analyzes the data to derive relevant EEG features, such as DE and PSD features, based on the specified feature types and frequency bands provided by the user. Sample segmentation organizes the data into user-defined sample lengths. Ultimately, the processed data is integrated into the format of (session, subject, trial, sample). This functionality is illustrated in the usage examples of LibEER for data loading. A data loader usage example is presented in Fig. 4.
@@ IV-2 Data split
To accommodate various splitting methods, the get_split_index() function further divides the data based on the specified formats and labels. The splitting methods available in LibEER include train-test splitting, which partitions the data into training and testing sets according to a predetermined ratio;train-val-test splitting, which divides the data into training, validation, and testing sets while maintaining balanced label distribution;and cross-validation methods, which segment the data into n folds, ensuring that each fold is utilized as the testing set once. The following example (see Fig. 5) illustrates how to employ LibEER for data splitting.
Figure 6: The example show how to create and train a model
@@
R2G-STNN SEED B, R, DE, 9s dependent 3 : 2 ACC 93.38$\pm$5.96 84.11$\pm$8.47 9.27$\downarrow$
MS-MDA SEED B, DE, 1s cross 14 : 1 ACC 89.63 93.97 4.34$\uparrow$
@@
HSLT DEAP-V B, R, 6s cross 14 : 1 ACC, F1 66.51 69.18 2.67$\uparrow$
DEAP-A B, R, 6s cross 14 : 1 ACC, F1 65.75 68.81 3.06$\uparrow$
DEAP-VA B, R, 6s cross 14 : 1 ACC, F1 56.93 49.57 7.36$\downarrow$
@@ IV-3 Model training and evaluation
Model training and evaluation are essential processes responsible for training the designated model and assessing its performance. LibEER constructs models based on the selected architecture and parameters, subsequently training them according to the specified training configurations. In scenarios where a validation set is not available, LibEER identifies the optimal performance on the testing set during training as the final outcome for that training iteration. Conversely, when a validation set is utilized, LibEER selects the network weights that yield the best performance on the validation set for evaluation against the testing set, thereby determining the final result of the training round.
The ultimate outcomes of the model will be reported according to the performance metrics specified by the user. The following example in Fig. 6 illustrates the process of creating and training a model within LibEER, along with the corresponding evaluation metrics.
@@ V Experimental Results
In this section, we thoroughly validate the importance of LibEER and our benchmark through extensive experimentation. Initially, we utilize LibEER to replicate the selected models as closely as possible following the experimental configurations detailed in the original publications. Subsequently, we employ LibEER to compare the performance of all models based on the established benchmark. Lastly, we conduct ablation studies to address specific issues identified during the experimental phase.
@@
DBN HSLT EEGNet CDCN TSception ACRNN DGCNN RGNN GCBNet GCBNet_BLS
SEED ACC 75.08 71.88 64.83 58.81 68.23 64.01 49.71 82.55 76.55 80.56 76.64
(19.73) (19.02) (20.47) (16.22) (20.35) (16.44) (13.15) (15.61) (16.92) (16.98) (17.44)
F1 70.82 67.39 58.82 54.41 63.76 60.53 45.78 79.89 72.52 77.29 72.52
(23.51) (22.81) (23.36) (17.59) (24.49) (18.51) (14.18) (18.93) (20.08) (20.92) (21.20)
SEED-IV ACC 47.80 45.56 40.28 29.89 52.26 36.06 29.01 52.39 45.40 53.28 53.51
(23.03) (21.19) (23.80) (13.53) (21.97) (15.12) (7.10) (24.32) (22.90) (21.05) (22.45)
F1 40.17 37.61 30.92 26.59 45.26 32.77 19.80 45.94 38.24 46.26 46.91
(21.68) (20.68) (24.47) (13.58) (23.00) (15.08) (5.42) (24.17) (23.09) (22.27) (22.46)
HCI-V ACC 64.83 62.03 64.00 61.15 60.48 61.12 60.51 67.83 64.86 66.84 69.60
(25.95) (24.90) (11.40) (16.76) (21.90) (15.52) (16.89) (22.40) (17.36) (21.42) (22.09)
F1 55.99 52.84 55.77 50.35 51.73 50.51 49.39 54.78 50.41 54.61 57.78
(27.80) (27.04) (13.12) (17.28) (23.11) (16.69) (15.70) (26.69) (20.34) (25.08) (27.18)
HCI-A ACC 63.61 68.51 67.74 67.42 71.82 68.26 66.26 67.29 70.96 64.89 69.60
(23.44) (21.72) (17.22) (21.71) (21.72) (23.10) (22.69) (27.73) (19.79) (27.12) (22.09)
F1 50.99 57.18 58.42 54.50 62.89 56.29 55.17 58.04 57.66 57.43 57.78
(24.89) (26.63) (19.50) (20.05) (24.99) (23.79) (23.20) (31.41) (25.39) (29.50) (27.18)
HCI-VA ACC 46.29 44.38 46.99 38.32 52.00 40.00 41.00 53.06 49.46 49.15 49.24
(28.42) (26.78) (20.76) (19.51) (26.05) (20.60) (21.58) (24.44) (23.20) (26.16) (27.97)
F1 33.91 31.96 34.76 24.56 38.04 27.19 27.10 39.75 35.97 36.49 36.68
(27.65) (27.87) (19.69) (13.71) (26.88) (13.65) (15.13) (26.01) (23.84) (27.58) (27.25)
DEAP-V ACC 54.17 56.08 56.20 51.50 57.71 51.52 53.52 56.07 55.90 56.49 57.02
(18.67) (17.38) (18.14) (11.57) (14.72) (9.54) (9.29) (17.15) (16.24) (18.17) (15.07)
F1 49.73 48.61 48.21 47.85 53.41 47.33 48.31 49.08 47.25 50.36 51.40
(18.92) (19.33) (18.73) (11.70) (15.50) (9.58) (7.77) (17.50) (17.55) (19.57) (17.25)
DEAP-A ACC 63.49 64.60 59.74 61.30 63.37 57.49 61.83 62.68 66.09 65.95 61.07
(16.72) (19.42) (18.82) (15.88) (14.18) (11.86) (14.32) (19.66) (13.91) (17.61) (16.56)
F1 53.31 52.61 50.10 53.26 53.94 50.75 49.68 53.94 49.27 55.34 50.43
(14.39) (19.85) (18.08) (13.05) (13.76) (11.30) (9.20) (20.10) (12.86) (17.86) (16.36)
DEAP-VA ACC 37.32 39.50 43.34 39.41 38.08 35.68 38.20 41.86 44.53 38.80 37.51
(17.24) (13.99) (14.49) (11.53) (15.52) (12.08) (11.34) (11.57) (14.35) (16.23) (15.29)
F1 25.55 24.88 23.47 29.19 28.90 26.75 21.05 29.12 25.88 27.89 27.00
(14.23) (10.79) (14.21) (10.21) (13.15) (8.30) (6.13) (10.92) (12.08) (15.94) (13.52)
@@
TABLE VI: The Mean Accuracies and F1 scores using the proposed benchmark for Cross-Subject EER Experiment. The top two methods in each scenario are highlighted using bold and underlined formatting.
Method SVM DNN Trans CNN RNN GNN

### v3

@@ III-D Datasets
Research in the EER field exhibits variability in dataset selection and often lacks clarity regarding the specific data segments utilized. This study evaluates the performance of popular models using six datasets, which are selected by comprehensively considering factors such as their frequency of use, data scale, and data quality. They will be introduced along with the relevant data segments, as summarized in Table III.
@@ III-D3 DEAP
The DEAP[27] dataset consists of 32 subjects (16 males, 16 females) who participated in one session containing 40 trials. Each trial involved a 1-minute music video designed to evoke specific emotions, followed by participant ratings for arousal, valence, dominance, and liking on a 1-9 scale. EEG data were recorded using a 32-channel cap at 512 Hz, and the dataset includes both raw and downsampled (128 Hz) versions. All samples are used for both subject-dependent and cross-subject tasks. In the subsequent experiments, we used the DEAP dataset with the valence labels, the DEAP dataset with the arousal labels, and the DEAP dataset with both labels, denoted as DEAP-V, DEAP-A, and DEAP-VA, respectively.
@@ III-D4 MAHNOB-HCI
The MAHNOB-HCI[29] dataset includes 30 subjects (13 males, 17 females), though EEG data are completely missing for 2 subjects and partially missing for 3. Each subject participated in a single session with 20 trials, where videos lasting 34.9 to 117 seconds evoked emotions. After each video, participants rated arousal, valence, dominance, and predictability on a 1-9 scale. EEG data were recorded using a 32-channel system at 512 Hz, and the dataset includes both raw and downsampled (128 Hz) versions. Both complete and partial samples are used for subject-dependent and cross-subject tasks. In the subsequent experiments, we used the MAHNOB-HCI dataset with the valence labels, the MAHNOB-HCI dataset with the arousal labels, and the MAHNOB-HCI dataset with both labels, denoted as HCI-V, HCI-A, and HCI-VA, respectively.
III-D5 SEED-V
The SEED-V[31] dataset contains 16 subjects (6 males, 10 females) who each participated in three experimental sessions. Each session included 15 emotion-inducing movie clips (3 clips per emotion category: happy, sad, neutral, fear, and disgust). EEG signals were recorded using a 62-channel cap following the 10-20 system, accompanied by simultaneous eye movement recordings. The dataset provides differential entropy (DE) features extracted from five frequency bands (delta, theta, alpha, beta, and gamma) for EEG signals. Data from all three sessions are used for subject-dependent and cross-subject tasks.
III-D6 MPED
The MPED[32] dataset is a comprehensive multimodal resource for discrete emotion recognition, comprising physiological recordings from 23 Chinese participants (10 males, 13 females). During emotion elicitation experiments using 28 standardized video stimuli representing seven discrete emotions (joy, funny, anger, sadness, disgust, fear, neutrality), 62-channel EEG signals were acquired via an ESI NeuroScan System at 1000 Hz sampling rate following the international 10-20 system. All data are used for both subject-dependent and cross-subject tasks.
@@ III-E Evaluation Methods and Metrics
where TP (True Positives), TN (True Negatives), FP (False Positives), and FN (False Negatives) are the metrics used to calculate accuracy and F1 score. Besides, the random seed is fixed at 2024 to ensure the reproducibility of the results.
@@ IV LibEER Toolkit Framework
Figure 4: The framework and data pipeline of LibEER.
LibEER offers a user-friendly and equitable platform for EER implemented using PyTorch. It comprises three primary modules: the data loader, data splitting, and model training and evaluation. The data loader module standardizes the data formats across various datasets and provides a range of preprocessing techniques. The data splitting module delivers a cohesive solution applicable to different experimental tasks and partitioning methods. The model training and evaluation module presents a standardized and extensible framework for conducting model training and evaluation. Fig. 4 illustrates the framework of LibEER. Below, we provide detailed explanations for each module.
@@ IV-1 Data loader
Additionally, LibEER provides the preprocess() function, which allows for the preprocessing of data based on user-defined settings. The preprocess() function operates on a trial basis and encompasses three primary steps: noise removal, feature extraction, and sample segmentation. When utilizing raw data directly from the dataset, only sample segmentation is executed. Noise removal aims to eliminate electromyography (EMG) noise, electrooculography (EOG) artifacts, and other forms of interference from the EEG signals. The feature extraction process analyzes the data to derive relevant EEG features, such as DE and PSD features, based on the specified feature types and frequency bands provided by the user. Sample segmentation organizes the data into user-defined sample lengths. Ultimately, the processed data is integrated into the format of (session, subject, trial, sample). An example demonstrating how to use the data loader is provided in our GitHub project.
@@ IV-2 Data split
To accommodate various splitting methods, the get_split_index() function further divides the data based on the specified formats and labels. The splitting methods available in LibEER include train-test splitting, which partitions the data into training and testing sets according to a predetermined ratio;train-val-test splitting, which divides the data into training, validation, and testing sets while maintaining balanced label distribution;and cross-validation methods, which segment the data into n folds, ensuring that each fold is utilized as the testing set once. An example illustrating how to perform data splitting with LibEER is provided in our GitHub project.
@@
TSception DEAP-V B, R, 4s dependent 9 : 1 ACC 61.57$\pm$11.04 61.89$\pm$12.16 0.32$\uparrow$
DEAP-A B, R, 4s dependent 9 : 1 ACC 59.14$\pm$7.60 60.02$\pm$8.84 0.88$\uparrow$
@@
SEED B, R, DE, 9s cross 14 : 1 ACC 83.28$\pm$9.60 80.36$\pm$8.25 2.92$\downarrow$
R2G-STNN SEED B, R, DE, 9s dependent 3 : 2 ACC 93.38$\pm$5.96 88.71$\pm$8.85 4.67$\downarrow$
SEED B, R, DE, 9s cross 14 : 1 ACC 84.16$\pm$7.63 78.35$\pm$9.46 5.81$\downarrow$
NSAL-DGAT SEED B, R, DE, 1s dependent 3 : 2 ACC 96.53$\pm$9.10 92.27$\pm$10.09 4.26$\downarrow$
SEED B, R, DE, 1s cross 14 : 1 ACC 95.69$\pm$5.24 90.19$\pm$7.32 5.50$\downarrow$
DAN SEED B, R, DE, 1s cross 14 : 1 ACC 83.81$\pm$8.56 76.29$\pm$6.60 7.52$\downarrow$
DANN SEED B, R, DE, 1s cross 14 : 1 ACC —— 77.40$\pm$5.86 ——
MS-MDA SEED B, DE, 1s cross 14 : 1 ACC 89.63$\pm$6.79 89.65$\pm$7.24 0.02$\uparrow$
PR-PL SEED B, R, DE, 1s dependent 3 : 2 ACC 94.84$\pm$9.16 93.2$\pm$6.66 1.64$\downarrow$
SEED B, R, DE, 1s cross 14 : 1 ACC 94.84$\pm$9.16 89.19$\pm$5.34 5.65$\downarrow$
HSLT DEAP-V B, R, PSD, 6s cross 14 : 1 ACC, F1 66.51 69.18 2.67$\uparrow$
DEAP-A B, R, PSD, 6s cross 31 : 1 ACC, F1 65.75 68.81 3.06$\uparrow$
DEAP-VA B, R, PSD, 6s cross 31 : 1 ACC, F1 56.93 49.57 7.36$\downarrow$
@@ IV-3 Model training and evaluation
Model training and evaluation are essential processes responsible for training the designated model and assessing its performance. LibEER constructs models based on the selected architecture and parameters, subsequently training them according to the specified training configurations. In scenarios where a validation set is not available, LibEER identifies the optimal performance on the testing set during training as the final outcome for that training iteration. Conversely, when a validation set is utilized, LibEER selects the network weights that yield the best performance on the validation set for evaluation against the testing set, thereby determining the final result.
The ultimate outcomes of the model will be reported according to the performance metrics specified by the user. An example of creating, training, and evaluating a model within LibEER is provided in our GitHub project.
@@ V Experimental Results
In this section, we thoroughly validate the importance of LibEER and our benchmark through extensive experimentation. Initially, we utilize LibEER to replicate the selected models as closely as possible following the experimental configurations detailed in the original publications. Subsequently, we employ LibEER to compare the performance of all models based on the established benchmark. Lastly, we investigate the impact of several key experimental settings on model performance through comparative experiments.
@@
EEGNet CDCN TSception ACRNN BiDANN R2G-STNN DGCNN GCBNet GCBNet _BLS RGNN NSAL-DGAT DBN DAN DANN PR-PL HSLT
SEED ACC 75.08 58.81 68.23 64.01 49.71 78.37 78.88 82.55 80.56 76.64 76.55 78.62 71.88 60.35 62.00 76.54 64.83
(19.73) (16.22) (20.35) (16.44) (13.15) (20.36) (17.34) (15.61) (16.98) (17.44) (16.92) (17.48) (19.02) (17.03) (21.95) (18.80) (20.47)
F1 70.82 54.41 63.76 60.53 45.78 75.91 73.72 79.89 77.29 72.52 72.52 74.77 67.39 53.12 53.58 72.19 58.82
(23.51) (17.59) (24.49) (18.51) (14.18) (24.60) (21.20) (18.93) (20.92) (21.20) (20.08) (20.95) (22.81) (19.70) (26.90) (22.85) (23.36)
SEED-IV ACC 47.80 29.89 52.26 36.06 29.01 51.82 53.04 52.39 53.28 53.51 45.40 53.48 45.56 44.78 46.49 54.59 40.28
(23.03) (13.53) (21.97) (15.12) (7.10) (23.71) (23.62) (24.32) (21.05) (22.45) (22.90) (21.21) (21.19) (23.84) (24.67) (22.13) (23.80)
F1 40.17 26.59 45.26 32.77 19.80 44.65 46.48 45.94 46.26 46.91 38.24 47.09 37.61 37.18 37.76 48.24 30.92
(21.68) (13.58) (23.00) (15.08) (5.42) (23.72) (24.76) (24.17) (22.27) (22.46) (23.09) (20.92) (20.68) (22.60) (23.89) (22.32) (24.47)
HCI-V ACC 64.83 61.15 60.48 61.12 60.51 68.50 63.10 67.83 66.84 70.60 64.86 65.17 62.03 63.55 58.94 61.18 64.00
(25.95) (16.76) (21.90) (15.52) (16.89) (24.18) (23.41) (22.40) (21.42) (21.05) (17.36) (20.52) (24.90) (22.19) (19.16) (12.19) (11.40)
F1 55.99 50.35 51.73 50.51 49.39 58.74 53.30 54.78 54.61 58.96 50.41 58.85 52.84 48.68 44.96 41.16 55.77
(27.80) (17.28) (23.11) (16.69) (15.70) (27.88) (25.38) (26.69) (25.08) (27.45) (20.34) (23.09) (27.04) (23.25) (21.05) (11.22) (13.12)
HCI-A ACC 63.61 67.42 71.82 68.26 66.26 67.62 65.72 67.29 64.89 69.60 70.96 65.25 68.51 61.67 64.54 64.57 67.74
(23.44) (21.71) (21.72) (23.10) (22.69) (24.58) (25.05) (27.73) (27.12) (22.09) (19.79) (26.37) (21.72) (25.08) (19.95) (23.95) (17.22)
F1 50.99 54.50 62.89 56.29 55.17 57.77 54.82 58.04 57.43 57.78 57.66 53.40 57.18 48.25 50.16 50.50 58.42
(24.89) (20.05) (24.99) (23.79) (23.20) (28.86) (24.52) (31.41) (29.50) (27.18) (25.39) (27.67) (26.63) (27.22) (23.81) (26.39) (19.50)
HCI-VA ACC 46.29 38.32 52.00 40.00 41.00 51.31 48.41 53.06 49.15 49.24 49.46 53.35 44.38 41.88 40.68 44.27 46.99
(28.42) (19.51) (26.05) (20.60) (21.58) (31.24) (25.36) (24.44) (26.16) (27.97) (23.20) (27.07) (26.78) (23.17) (22.11) (23.08) (20.76)
F1 33.91 24.56 38.04 27.19 27.10 37.62 36.10 39.75 36.49 36.68 35.97 38.42 31.96 26.62 24.72 29.90 34.76
(27.65) (13.71) (26.88) (13.65) (15.13) (27.81) (22.71) (26.01) (27.58) (27.25) (23.84) (27.48) (27.87) (17.51) (16.79) (20.68) (19.69)
DEAP-V ACC 54.17 51.50 57.71 51.52 53.52 55.21 55.96 56.07 56.49 57.02 55.90 52.21 56.08 57.89 57.70 59.23 56.20
(18.67) (11.57) (14.72) (9.54) (9.29) (19.70) (18.30) (17.15) (18.17) (15.07) (16.24) (16.19) (17.38) (19.18) (19.93) (14.69) (18.14)
F1 49.73 47.85 53.41 47.33 48.31 47.23 49.84 49.08 50.36 51.40 47.25 48.87 48.61 50.97 50.23 47.42 48.21
(18.92) (11.70) (15.50) (9.58) (7.77) (18.05) (17.77) (17.50) (19.57) (17.25) (17.55) (18.34) (19.33) (18.88) (20.97) (17.67) (18.73)
DEAP-A ACC 63.49 61.30 63.37 57.49 61.83 63.57 63.78 62.68 65.95 61.07 66.09 60.31 64.60 60.51 62.38 62.15 59.74
(16.72) (15.88) (14.18) (11.86) (14.32) (20.33) (18.92) (19.66) (17.61) (16.56) (13.91) (17.29) (19.42) (17.32) (16.78) (15.90) (18.82)
F1 53.31 53.26 53.94 50.75 49.68 52.93 48.68 53.94 55.34 50.43 49.27 50.21 52.61 50.96 48.57 53.22 50.10
(14.39) (13.05) (13.76) (11.30) (9.20) (20.26) (17.91) (20.10) (17.86) (16.36) (12.86) (18.91) (19.85) (16.53) (16.96) (17.99) (18.08)
DEAP-VA ACC 37.32 39.41 38.08 35.68 38.20 39.30 40.75 41.86 38.80 37.51 44.53 38.78 39.50 38.07 41.71 42.58 43.34
(17.24) (11.53) (15.52) (12.08) (11.34) (15.66) (15.79) (11.57) (16.23) (15.29) (14.35) (15.64) (13.99) (14.12) (15.51) (13.01) (14.49)
F1 25.55 29.19 28.90 26.75 21.05 27.32 28.63 29.12 27.89 27.00 25.88 25.75 24.88 25.91 28.88 27.63 23.47
(14.23) (10.21) (13.15) (8.30) (6.13) (13.23) (13.01) (10.92) (15.94) (13.52) (12.08) (11.92) (10.79) (10.98) (15.14) (11.43) (14.21)
SEED-V ACC 34.88 36.02 65.48 31.45 34.10 49.17 47.21 57.59 60.68 63.71 55.31 52.37 28.64 22.37 27.77 49.75 47.17
(18.42) (10.15) (11.49) (5.12) (7.21) (17.86) (19.22) (20.36) (18.42) (17.03) (12.88) (19.83) (15.62) (4.31) (17.64) (16.01) (16.82)
F1 29.83 26.77 49.51 20.05 26.29 36.82 38.02 48.59 50.13 51.31 46.55 46.25 22.42 7.54 20.02 39.25 31.61
(16.75) (8.11) (12.34) (2.59) (4.43) (14.18) (16.74) (20.88) (19.81) (18.91) (22.17) (19.34) (15.51) (2.16) (17.58) (15.51) (12.85)
MPED ACC 21.49 18.43 20.89 19.01 18.88 25.75 24.43 22.93 22.83 22.65 20.72 22.19 22.22 19.47 21.24 20.40 18.47
(7.33) (8.99) (5.48) (6.77) (9.97) (7.97) (9.28) (6.98) (8.30) (8.06) (5.34) (8.23) (5.17) (7.38) (7.52) (7.48) (4.28)
F1 17.97 17.80 17.00 12.63 13.27 21.26 21.03 19.10 18.88 19.00 19.06 18.67 16.76 14.83 17.50 17.86 16.54
(5.97) (6.58) (4.74) (6.80) (6.04) (6.81) (7.64) (6.28) (8.00) (6.70) (4.88) (7.45) (4.42) (7.39) (9.52) (6.61) (5.96)
@@
TABLE VI: The mean accuracies and F1 scores (and standard deviations) using the proposed benchmark for cross-subject EER experiment. The top two methods in each scenario are highlighted using bold and underlined formatting.
Method SVM CNN RNN GNN DNN Trans.